OpenStudy (anonymous):

(-3/4)(-7/15)(-8/21)(-5/9)

OpenStudy (texaschic101):

- 3/4 x - 7/15 = 21/60 = 7/20 (multiply first two) - 8/21 x - 5/9 = 40/189 (multiply last two) now multiply the answers you got ... 7/20 x 40/189 = 14/189 = 2/27 or... -3/4x - 7/15 = 7/20 (multiply first two) 7/20 x - 8/21 = -14/105 (multiply answer by 3rd number) -14/105 x - 5/9 = 14/189 = 2/27 (multiply answer by 4th number) either way, you will get the same answer... 2/27

OpenStudy (anonymous):

Thanksssssssss

OpenStudy (texaschic101):

no problem...thats what I am here for :)